OpenStudy (anonymous):

I need help with this equation. x^2+12x+36=25 I got -11, -1 I'm not sure if I placed the neg/pos signs incorrectly. . but, it's not an answer choice.

OpenStudy (haseeb96):

let me check that

OpenStudy (haseeb96):

okay would u wait 5 mins i willl give u answer

OpenStudy (phi):

we can test your numbers. For x=-1 x^2 = -1*-1 = 1 12 x is -12 and 1-12+36 simplifies to 36-11= 25 which matches now let's test -11

OpenStudy (phi):

-11*-11 = +121 12x is 12*-11 = -132 121-132+36 is 25 that also works

OpenStudy (haseeb96):

x^2+12x+36 =25 x^2 +12x +36-25=0 x^2 +12x +11=0 use the factorizing method x^2 +x +11x +11=0 x(x+1) +11(x+1) =0 (x+11)(x+1)=0 either' x+11=0 or x+1=0 x=-11 or x=-1 your answer is absolutely correct i m sure 100 %

OpenStudy (haseeb96):

@-EtherealEbullience match your answer with mine

OpenStudy (anonymous):

. . . Nevermind, I got it. xD

OpenStudy (phi):

when you solved this problem , you would do the following: x^2+12x+36=25 add -25 to both sides x^2+12x+36 -25 =25 - 25 and simplify to get x^2 +12x +11 = 0 It is in standard form. this factors: the 11 is 1*11 and the 1+11 is 12. This means we use 1 and 11 and write the factors (x+11)(x+1)=0 to solve, either x+11 =0 or x+1 = 0 in the first case x+11=0 add -11 to both sides. we get x= -11 and for the second case x+1=0, add -1 to both sides. we get x=-1

OpenStudy (anonymous):

Thank you, it was, I overlooked it.